OpenStudy (anonymous):

solve the equation -70=d/10

OpenStudy (anonymous):

Multiply both sides by 10.

OpenStudy (anonymous):

-70*10=? d/10*10=?

OpenStudy (anonymous):

The multiplication undoes the division of d/10, but you need to do it to both sides, because there is an equal sign.

OpenStudy (anonymous):

\[-70\times10=-700\]\[d/10\times10=d\]

OpenStudy (anonymous):

So, d=?

OpenStudy (anonymous):

-700

OpenStudy (anonymous):

Should give me a medal man.

OpenStudy (anonymous):

i can :) @OrangeMaster here u go